OpenStudy (anonymous):

What is the x- intercept of 5x - 2y = 10? (2, 0) (-2, 0) (5, 0) (-5, 0)

hero (hero):

5x - 2y = 10 To find the x intercept, set y = 0, then solve for x as follows: 5x - 2(0) = 10 5x - 0 = 10 5x = 10 x = 10/5 x = 2 Therefore, the x-intercept is (2,0)

OpenStudy (anonymous):

thanks